# Nonlinear spin and orbital Rashba-Edelstein effects induced by a femtosecond laser pulse: Simulations for Au(001)

### v1

@@ Abstract
Condensed matter physics, ultrafast electron dynamics, Nonlinear Rashba-Edelstein effect, Hall effect, tight-binding model, numerical simulations
@@ I Introduction
Rashba-type spin-orbit coupling (RSOC) manifests itself in a number of effects at surfaces or interfaces [1, 2]. A prominent example is the spin-momentum locking of surface states, for example, in the L-gap surface state in Au(111) [3], and the accompanying spin splitting, for example, in the surface alloy Bi/Ag(111) [4]. Concerning nonequilibrium situations, the (Rashba-)Edelstein effect is a clear signature of RSOC [5, 6, 7, 8, 9]: an electric current flowing parallel to a surface produces a homogeneous spin polarization in the surface layers that is in-plane and perpendicular to the current’s direction. This spin Edelstein effect (SEE) is complemented by the orbital Edelstein effect (OEE), in which the current results in an orbital polarization [10, 11, 12, 13, 14, 9, 15].
Both the spin and the orbital Edelstein effect are well understood for the steady state in which the applied electric field is time-independent. For ultrafast electron dynamics, results obtained for the steady state cannot be easily transferred into the time domain. An example is the ultrafast orbital Hall effect generated by a femtosecond laser pulse, which exhibits similarities but also striking differences in comparison to its steady-state counterpart [16].
@@
Figure 1: Sketch of the ultrafast spin and orbital Edelstein effect in a two-dimensional system. (a) A linearly polarized laser pulse with electric field $E$ along the $+x$-direction (violet) impinges perpendicular to the sample (gray; located in the $xy$ plane). $O_{y}$ (red) is the laser-induced in-plane angular momentum $O$ (either spin or orbital angular momentum) along the $+y$-direction. (b) Same as (a) but half a period later: $E$ and $O_{y}$ point along the $-x$- and the $-y$-direction, respectively.
@@
Ultrafast phenomena can be delineated on the level of time-dependent density functional theory (TDDFT), a very accurate description [17, 18, 19]. Being computationally demanding, TDDFT simulations are usually restricted to small sample sizes and short time spans. By contrast, model calculations, which require less computational effort, may not capture important details of the electronic structure, in particular, if highly excited electronic states are involved in the dynamics. Our computational approach evolve [20, 21, 22, 23, 16] bridges the gap between TDDFT and model calculations: based on a tight-binding method in real space, it allows treating comparably large samples and long time spans, and it represents the electronic structure reasonably well. evolve has proven its suitability in a number of investigations. For example, a backflow mechanism in Co/Cu heterostructures is faithfully reproduced [20]. The tight-binding description imposes two restrictions: the discrete representation in terms of sites and atomic orbitals imposes, first, that spatial resolution is limited to atomic sites and, second, the hopping matrix elements in the Hamilton operator introduce a lower limit of the temporal resolution (in the order of a few tens of attoseconds).
@@ II Theoretical aspects
We briefly introduce the main ideas of our approach to ultrafast electron dynamics. The computations have been performed with the computer code evolve, which is being developed in our group (for details, see Refs. 20, 21, 22, 23, 16).
The samples are Au(001) monolayers. The free-standing films form a square lattice, with Cartesian axes chosen as $x\equiv[110]$, $y\equiv[\bar{1}10]$, and $z\equiv[001]$. The samples are approximated by real-space clusters with dimension $11\times 1$ 11 atomic sites, with periodic boundary conditions applied in the $x$- and in $y$-directions (i. e. closed circuit geometry). Au lends itself to a study of ultrafast SEE and OEE since it exhibits a sizable spin Hall conductivity [24, 25, 26] and a moderate Rashba spin-splitting [27, 28, 29, 30].
The electronic structure of the samples is described by a semi-empirical real-space tight-binding (TB) approach [31]. The Hamiltonian $\hat{H}_{0}$ is of Slater-Koster-type with parameters taken from Ref. 32. It includes both the atomic spin-orbit coupling [33] and the RSOC for gold [34]. The latter mimics the effect of a substrate which induces a potential gradient perpendicular to the film; for details, see Appendix B.
The electronic system described by $\hat{H}_{0}$ is excited by a femtosecond laser pulse [35], whose electric field oscillates within the sample along the $x$-axis. This field has a sinusoidal carrier wave with an energy of $\hbar\omega=3.0\,\mathrm{eV}$ (equivalent to a period $T$ of about $1.38\,\mathrm{fs}$) and a Lorentzian envelope of $10\,\mathrm{fs}$ width and center at $t=0\,\mathrm{fs}$. The photon energy was chosen to allow excitation of the occupied $d$-states (centered about $3\,\mathrm{eV}$ below the chemical potential) into the unoccupied $p$-orbitals. The laser’s amplitude is chosen to obtain a fluence of about $3.2\,\mathrm{mJ\,cm^{-2}}$ and $0.15$ excited electrons per site (at the laser’s maximum; about $0.05$ thereafter). The geometry of the entire setup (sample and laser) dictates that only certain components of the SAM and the OAM are produced by the incident radiation (see Refs. 36, 22, 23, 16, as well as the symmetry analysis in Appendix A).
@@
for the one-particle density matrix $\hat{\rho}(t)$. The latter is either expressed in a site-orbital-spin basis or in the eigenstate basis of $\hat{H}_{0}$. The time-dependent Hamiltonian $\hat{H}(t)$ supplements $\hat{H}_{0}$ by the rapidly oscillating laser radiation. The latter is taken into account by a unitary transformation [35] which includes dipole transitions in first order in the electric field.
@@
at site $k$ ($\hat{\boldsymbol{\sigma}}$ vector of Pauli matrices, $\hat{\boldsymbol{L}}=(\hat{L}_{x},\hat{L}_{y},\hat{L}_{z})$ vector of OAM operators). Orbital momenta are calculated in the atomic center approximation (ACA) [11, 37, 16] and are brought about by hybridization of orbitals [38, 39].
The electric field of the laser does not only change the above quantities locally but also results in respective currents [40, 22, 16]. The probability current from site $l$ to site $k$ is calculated from
@@
$\rho_{lk}$ and $h_{kl}$ are off-diagonal blocks of the density matrix and of the Hamiltonian matrix in the site representation, respectively. Transport of the $\mu$-th momentum component from site $k$ to site $l$ is quantified by [23, 16]
@@ III Results and discussion
A photo excitation of an electron system produces, in general, SAM and OAM, an effect well-known from photoelectron emission spectroscopy [41, 42, 43, 44, 36, 45, 46]. We performed a symmetry analysis in order to identify the RSOC-induced SAM and OAM components.
@@ III.1 Laser-induced nonlinear Edelstein effects
The photo-induced SAM and OAM components do not follow the laser intensity; their envelope differs from a Lorentzian shape (cf. Figure 2). The induced moments do not scale with the laser’s electric field amplitude and thus correspond to a nonlinear Edelstein effect. Supporting this viewpoint, additional simulations confirm that the spin and orbital responses in our system exhibit clear nonlinear dependence on the amplitude of the driving electric field. Accordingly, we interpret the laser-induced signals in Figure 2 as manifestations of nonlinear SEE and OEE, respectively.
In the data, pronounced maxima appear at about $7.5\,\mathrm{fs}$ (panels b and c), which is clearly after the laser’s maximum at $0\,\mathrm{fs}$. Moreover, both signals exhibit distinct beating patterns after the laser pulse has faded away. In a recent publication, we related the response to the laser field (orbital Hall currents in that case [16]) to the classical analog of a driven oscillator. This suggests relating the beating patterns studied here to those of a driven pendulum at low amplitudes; for the latter, beating is observed between the transient and steady-state response [47].
@@
In the present simulations, the coupling of the electron system to a bosonic heat bath via Lindblad operators has been neglected [20, 21], as our primary focus is on the fundamental processes that are induced by the laser excitation. Nonetheless, we performed additional calculations that include this coupling. As expected, the key processes discussed here still appear. The magnitude and overall features (e.g., rapid oscillations and beating patterns) of the resulting signals are slightly modified within the time range shown in the figures. Consistent with expectations, the presence of the heat bath leads to a reduction in signal amplitude following the pulse, occurring within a few hundred femtoseconds, depending on the coupling parameters.
Within the evolve framework, only the intrinsic contributions to the Edelstein effects are considered, as the system is assumed to be idealfree from defects and thus lacking any extrinsic contributions. This naturally raises the question: why do Edelstein effects manifest on ultrafast timescales?
Since the laser-induced magnetic moments depend nonlinearly on the laser’s electric field, the intrinsic Edelstein effect is nonlinear and distinct from the linear extrinsic Edelstein effect that is typically discussed in a Rashba system. In the context of linear response theory, as described by the Kubo formalism, the emergence of a linear extrinsic Edelstein effect is closely tied to the symmetries of time reversal $\mathcal{T}$ and spatial inversion $\mathcal{I}$. A broken $\mathcal{I}$ symmetry is necessary for both extrinsic and intrinsic contributions, ensured here by the out-of-plane potential gradient (Appendix A). However in linear response, a nonvanishing intrinsic contribution also requires the breaking of $\mathcal{T}$ symmetry, even in a nonmagnetic system. Thus, the breaking of time-reversal symmetry becomes essential for the linear Edelstein effects, as detailed in Ref. 9.
@@
A nonlinear Edelstein effect  as we observe here, can occur independently of time-reversal symmetry. This was explicitly demonstrated by Vignale and Tokatly in the case of a noninteracting, ideal two-dimensional electron gas (2DEG) [48]. Laser-induced nonlinear Edelstein effects in nonmagnetic systems have previously been attributed to energy dissipation associated with photocurrents generated during laser excitation [49]. When the laser pulse interacts with the system, it injects energy and couples electronic states that were previously disconnected. This results in a redistribution of population among the energy levels: the density matrix becomes more complex, acquiring additional nonzero elements both on and off its diagonal.
@@ III.2 Laser-induced currents
The longitudinal $y$-polarized SAM current (panel b in Figure 3) exhibits a frequency that is twice as large as that of the laser pulse. This feature has been observed for the ultrafast orbital Hall effect [16]; it is explained in a “two-current model”: the sign of the laser-induced $y$-polarized SAM is strictly related to the sign of the laser’s electric field (confer Appendix A); this also holds for the probability current. In one half-period of the laser oscillation, positive $y$-polarized SAM is transported along the $+x$ direction. In the next half-period, negative $y$-polarized SAM is transported along $-x$; the latter is equivalent to the transport of positive $y$-polarized SAM along $+x$. This results in the doubled frequency of the longitudinal $y$-polarized SAM current. The same argument holds for the OAM current (panel c in Figure 3). Both SAM and OAM currents exhibit a beating pattern. However, in contrast to $\braket{s_{y}}(t)$ and $\braket{l_{y}}(t)$ (Figure 2), their average is not small.
@@
The transversal currents depicted in Figure 4 substantiate that SAM and OAM currents behave differently, which may be attributed to the very nature of SAM and OAM: SAM is an intrinsic feature of electrons, while OAM is motion-related and determined by the lattice geometry as well as by the hybridization of the (occupied) orbitals (cf. the ACA [11, 38, 39]).
Additionally, a frequency analysis can help to determine the origin of the currents. Specifically, RSOC-induced currents [$\braket{j_{x}^{O_{y}}}(t)$ and $\braket{j_{y}^{O_{x}}}(t)$] oscillate at even multiples of the laser frequency, whereas currents [$\braket{j_{x}}(t)$ and $\braket{j_{y}^{O_{z}}}(t)$] that exist without RSOC (but with atomic spin-orbit coupling) oscillate at odd multiples of $\omega$ (for details, see Appendix D).

### v2

@@ I Introduction
Rashba-type spin-orbit coupling (RSOC) manifests itself in a number of effects at surfaces or interfaces Bychkov and Rashba (1984); Bihlmayer et al. (2022). A prominent example is the spin-momentum locking of surface states, for example, in the L-gap surface state in Au(111) Henk et al. (2004), and the accompanying spin splitting, for example, in the surface alloy Bi/Ag(111) Ast et al. (2007). Concerning nonequilibrium situations, the (Rashba-)Edelstein effect is a clear signature of RSOC Aronov and Lyanda-Geller (1989); Edelstein (1990); Inoue et al. (2003); Gambardella and Miron (2011); Johansson (2024): an electric current flowing parallel to a surface produces a homogeneous spin polarization in the surface layers that is in-plane and perpendicular to the current’s direction. This spin Edelstein effect (SEE) is complemented by the orbital Edelstein effect (OEE), in which the current results in an orbital polarization Zhong et al. (2016); Go et al. (2017); Yoda et al. (2018); Salemi et al. (2019); Johansson et al. (2021); Johansson (2024); Göbel et al. (2025).
Both the spin and the orbital Edelstein effect are well understood for the steady state in which the applied electric field is time-independent. For ultrafast electron dynamics, results obtained for the steady state cannot be easily transferred into the time domain. An example is the ultrafast orbital Hall effect generated by a femtosecond laser pulse, which exhibits similarities but also striking differences in comparison to its steady-state counterpart Busch et al. (2024).
@@
Figure 1: Sketch of the ultrafast spin and orbital Edelstein effect in a two-dimensional system. (a) A linearly polarized laser pulse with electric field $E$ along the +x+x-direction (violet) impinges perpendicular to the sample (gray; located in the xyxy plane). $O_{y}$ (red) is the laser-induced in-plane angular momentum $O$ (either spin or orbital angular momentum) along the +y+y-direction. (b) Same as (a) but half a period later: $E$ and $O_{y}$ point along the $x-$ and the y-y-direction, respectively.
@@
Ultrafast phenomena can be delineated on the level of time-dependent density functional theory (TDDFT), a very accurate description Ullrich (2011); Krieger et al. (2015); Dewhurst et al. (2018). Being computationally demanding, TDDFT simulations are usually restricted to small sample sizes and short time spans. By contrast, model calculations, which require less computational effort, may not capture important details of the electronic structure, in particular, if highly excited electronic states are involved in the dynamics. Our computational approach evolve Töpler et al. (2021); Ziolkowski et al. (2023); Busch et al. (2023a, b, 2024) bridges the gap between TDDFT and model calculations: based on a tight-binding method in real space, it allows treating comparably large samples and long time spans, and it represents the electronic structure reasonably well. evolve has proven its suitability in a number of investigations. For example, a backflow mechanism in Co/Cu heterostructures is faithfully reproduced Töpler et al. (2021). The tight-binding description imposes two restrictions: the discrete representation in terms of sites and atomic orbitals imposes, first, that spatial resolution is limited to atomic sites and, second, the hopping matrix elements in the Hamilton operator introduce a lower limit of the temporal resolution (in the order of a few tens of attoseconds).
@@ II Theoretical aspects
We briefly introduce the main ideas of our approach to ultrafast electron dynamics. The computations have been performed with the computer code evolve, which is being developed in our group (for details, see Refs. Töpler et al., 2021; Ziolkowski et al., 2023; Busch et al., 2023a, b, 2024).
The samples are Au(001) monolayers. The free-standing films form a square lattice, with Cartesian axes chosen as $x\equiv[110]$, $y\equiv[\bar{1}10]$, and $z\equiv[001]$. The samples are approximated by real-space clusters with dimension $11\times 11$ atomic sites, with periodic boundary conditions applied in $x$- and in $y$-directions (i. e. closed circuit geometry). Au lends itself to a study of the ultrafast SEE and OEE since it exhibits a sizable spin Hall conductivity Seki et al. (2008); Isasa et al. (2015); Zou et al. (2016) and a moderate Rashba spin-splitting LaShell et al. (1996); Nicolay et al. (2001); Hoesch et al. (2004); Cercellier et al. (2006).
[…]
The electronic system described by $\hat{H}_{0}$ is excited by a femtosecond laser pulse Savasta and Girlanda (1995), whose electric field oscillates within the sample along the $x$-axis. This field has a sinusoidal carrier wave with an energy of $\hbar\omega=3.0\,\mathrm{eV}$ (equivalent to a period $T$ of about $1.38\,\mathrm{fs}$) and a Lorentzian envelope of $10\,\mathrm{fs}$ width and center at $t=0\,\mathrm{fs}$. The photon energy was chosen to allow excitation of the occupied $d$-states (centered about $3\,\mathrm{eV}$ below the chemical potential) into the unoccupied $p$-orbitals. The laser’s amplitude is chosen to obtain a fluence of about $3.2\,\mathrm{mJ\,cm^{-2}}$ and $0.15$ excited electrons per site (at the laser’s maximum; about $0.05$ thereafter). The geometry of the entire setup (sample and laser) dictates that only certain components of the SAM and the OAM are produced by the incident radiation (see Refs. Henk et al., 1996; Busch et al., 2023a, b, 2024, as well as the symmetry analysis in Appendix A).
@@
for the one-particle density matrix $\hat{\rho}(t)$. The latter is either expressed in a site-orbital-spin basis or in the eigenstate basis of $\hat{H}_{0}$. The time-dependent Hamiltonian $\hat{H}(t)$ supplements $\hat{H}_{0}$ by the rapidly oscillating laser radiation. The latter is taken into account by a unitary transformation Savasta and Girlanda (1995) which includes dipole transitions in first order in the electric field.
@@
at site $k$ ($\hat{\boldsymbol{\sigma}}$ vector of SAM operators, $\hat{\boldsymbol{L}}=(\hat{L}_{x},\hat{L}_{y},\hat{L}_{z})$ vector of OAM operators). Orbital momenta are calculated in the atomic center approximation (ACA) Go et al. (2017); Pezo et al. (2022); Busch et al. (2024) and are brought about by hybridization of orbitals Go et al. (2018, 2021).
The electric field of the laser does not only change the above quantities locally but also results in respective currents Mahan (2000); Busch et al. (2023a, 2024). The probability current from site $l$ to site $k$ is calculated from (in atomic units)
@@
$\rho_{lk}$ and $h_{kl}$ are off-diagonal blocks of the density matrix and of the Hamiltonian matrix in the site representation, respectively. Transport of the $\mu$-th momentum component from site $k$ to site $l$ is quantified by Busch et al. (2023b, 2024)
@@ III Results and discussion
A photo excitation of an electron system produces, in general, SAM and OAM, an effect well-known from photoelectron emission spectroscopy Tamura et al. (1987); Tamura and Feder (1991a, b); Henk and Feder (1994); Henk et al. (1996); Ünzelmann et al. (2020); Kobayashi et al. (2020). We performed a symmetry analysis in order to identify the RSOC-induced SAM and OAM components.
@@
In general, the OEE requires broken inversion symmetry Johansson (2024). However, unlike the SEE, it does not necessarily rely on spin–orbit interaction. In periodic systems, an orbital texture can arise, for instance, in chiral materials Yoda et al. (2018); Göbel et al. (2025), where inversion symmetry is inherently broken, and energy splitting of spin-degenerate bands is not required. Such orbital textures—often antisymmetric—can result from orbital hybridization Go et al. (2017, 2018) or be described using the modern framework, which also captures topological aspects Xiao et al. (2006); Shi et al. (2007).
In the case of the Au film studied here, the intrinsic crystal structure is inversion symmetric. However, inversion symmetry is effectively broken by the RSOC term in the Hamiltonian, which can emerge due to the influence of a substrate. This scenario is analogous to the conventional Rashba model, where inversion symmetry breaking gives rise to the SEE. In our system, however, excitation is induced by a laser pulse, and the dipole-transition matrix elements play a crucial role. This necessitates the inclusion of atomic SOC to facilitate the required orbital mixing. Consequently, both SEE and OEE appear in the system.
@@ III.1 Laser-induced nonlinear Edelstein effects
The photo-induced SAM and OAM components do not follow the laser intensity; their envelope differs from a Lorentzian shape (cf. Figure 2). The induced moments do not scale with the laser’s electric field amplitude and thus correspond to a nonlinear Edelstein effect Ye et al. (2024). Supporting this viewpoint, additional simulations confirm that the spin and orbital responses in our system exhibit clear nonlinear dependence on the amplitude of the driving electric field. Accordingly, we interpret the laser-induced signals in Figure 2 as manifestations of nonlinear SEE and OEE, respectively.
In the data, pronounced maxima appear at about $7.5\,\mathrm{fs}$ (panels b and c), which is clearly after the laser’s maximum at $0\,\mathrm{fs}$. Moreover, both signals exhibit distinct beating patterns after the laser pulse has faded away. In a recent publication, we related the time-resolved response to the laser field (orbital Hall currents in that case Busch et al. (2024)) to the classical dynamics analog of a driven damped oscillator. Such a model is able to qualitatively capture characteristic phase relations (notably the phase shift of $\pi/2$ between the orbital Hall currents and the driving field in Ref. Busch et al., 2024). This motivates comparing the beating patterns observed in Fig. 2 to those in classical systems, such as a driven pendulum at low amplitudes Cross (2021). For the latter, beating is observed between the transient and steady-state response. However, we stress that this analogy serves an illustrative purpose and should not be viewed as a rigorous explanation of the microscopic quantum dynamics.
Regarding the possible origins of the observed beating patterns—particularly whether multiphoton processes are involved—we note that the applied laser pulse is dominated by its carrier frequency [cf. Figure 6(a)]. In our computational framework, evolve, the laser’s electromagnetic field is treated classically, rendering the concept of photons—quantized excitations of the field—strictly speaking inapplicable. However, the simulations reveal a nonlinear dependency of the response amplitudes on the driving electric field. Such nonlinear behaviour is sometimes associated with multiphoton physics, although the classical treatment does not capture photon-based processes.
Consistent with earlier findings reported in Ref. Töpler et al., 2021, we observe that electronic states far above the chemical potential—at energies several times larger than the laser’s carrier energy—become transiently populated. This occurs through the continued evolution of the ‘excited’ density matrix $\rho(t)$, as governed by the von Neumann equation. Both excitation and de-excitation processes persist throughout the duration of the laser pulse. Since the induced population remains nearly unchanged after the pulse (see Figure 1 in Ref. Töpler et al., 2021), the beating patterns reflect coherent post-pulse dynamics of the excited electronic states rather than ongoing transitions.
@@
In the present simulations, the coupling of the electron system to a bosonic heat bath via Lindblad operators has been neglected Töpler et al. (2021); Ziolkowski et al. (2023), as our primary focus is on the fundamental processes that are induced by the laser excitation. Nonetheless, we performed additional calculations that include this coupling. As expected, the key processes discussed here still appear. The magnitude and overall features (e.g., rapid oscillations and beating patterns) of the resulting signals are slightly modified within the time range shown in the figures. Consistent with expectations, the presence of the heat bath leads to a reduction in signal amplitude following the pulse, occurring within a few hundred femtoseconds, depending on the coupling parameters.
Within the evolve framework, only the intrinsic contributions to the Edelstein effects are considered, as the system is assumed to be free from defects and thus lacking any extrinsic contributions. This naturally raises the question: why do Edelstein effects manifest on ultrafast timescales?
Since the laser-induced magnetic moments depend nonlinearly on the laser’s electric field, the intrinsic Edelstein effect is nonlinear and distinct from the linear extrinsic Edelstein effect that is typically discussed in a Rashba system. In the context of linear response theory, as described by the Kubo formalism, the emergence of a linear extrinsic Edelstein effect is closely tied to the symmetries of time reversal $\mathcal{T}$ and spatial inversion $\mathcal{I}$. A broken $\mathcal{I}$ symmetry is necessary for both extrinsic and intrinsic contributions, ensured here by the out-of-plane potential gradient (Appendix A). However in linear response, a nonvanishing intrinsic contribution also requires the breaking of $\mathcal{T}$ symmetry, even in a nonmagnetic system. Thus, within our model, the breaking of time-reversal symmetry would be essential for the linear intrinsic Edelstein effect, as detailed in Ref. Johansson, 2024.
@@
A nonlinear Edelstein effect  as we observe here, can occur independently of time-reversal symmetry. This was explicitly demonstrated by Vignale and Tokatly in the case of a noninteracting, ideal two-dimensional electron gas (2DEG) Vignale and Tokatly (2016). Laser-induced nonlinear Edelstein effects in nonmagnetic systems have previously been attributed to energy dissipation associated with photocurrents generated during laser excitation Xu et al. (2021). When the laser pulse interacts with the system, it injects energy and couples electronic states that were previously disconnected. This results in a redistribution of population among the energy levels: the density matrix becomes more complex, acquiring additional nonzero elements both on and off its diagonal.
@@ III.2 Laser-induced currents
The longitudinal $y$-polarized SAM current (panel b in Figure 3) exhibits a frequency that is twice as large as that of the laser pulse. This feature has been observed for the ultrafast orbital Hall effect Busch et al. (2024); it is explained in a “two-current model”: the sign of the laser-induced $y$-polarized SAM is strictly related to the sign of the laser’s electric field (confer Appendix A); this also holds for the probability current. In one half-period of the laser oscillation, positive $y$-polarized SAM is transported along the +x+x direction. In the next half-period, negative $y$-polarized SAM is transported along x-x; the latter is equivalent to the transport of positive $y$-polarized SAM along +x+x. This results in the doubled frequency of the longitudinal $y$-polarized SAM current. The same argument holds for the OAM current (panel c in Figure 3). Both SAM and OAM currents exhibit a beating pattern. However, in contrast to $\braket{s_{y}}(t)$ and $\braket{l_{y}}(t)$ (Figure 2), their average is not small.
@@
The transversal currents depicted in Figure 4 substantiate that SAM and OAM currents behave differently, which may be attributed to the very nature of SAM and OAM: SAM is an intrinsic feature of electrons, while OAM is motion-related and determined by the lattice geometry as well as by the hybridization of the (occupied) orbitals (cf. the ACA Go et al. (2017, 2018, 2021)).
Additionally, a frequency analysis can help to determine the origin of the currents. Specifically, RSOC-induced currents $\braket{j_{x}^{O_{y}}}(t)$ and $\braket{j_{y}^{O_{x}}}(t)$] oscillate at even multiples of the laser frequency, whereas currents $\braket{j_{x}}(t)$ and $\braket{j_{y}^{O_{z}}}(t)$] that exist without RSOC (but with atomic spin-orbit coupling) oscillate at odd multiples of $\omega$ (for details, see Appendix D).